OpenStudy (anonymous):

If two numbers are even, their sum is even. Which of these is a logically equivalent statement?

OpenStudy (anonymous):

If the sum of two numbers is not even, the numbers are not even. If two numbers are not even, their sum is even. If two numbers are not even, their sum is not even. If the sum of two numbers is not even, the numbers are even.

OpenStudy (anonymous):

'B'

OpenStudy (anonymous):

@JohnHanShanghai how?

OpenStudy (anonymous):

odd+odd=even

OpenStudy (anonymous):

can u give me an example?

OpenStudy (anonymous):

Use elimination here. The first statement is false because 2 + 3 = 5. Here, one number is even. Second one is true. 5 + 7 = 12. Third one is false. 3 + 7 = 10. Fourth one is also false. 3 + 4 = 7.

OpenStudy (anonymous):

So answer is the second one as @JohnHanShanghai has mentioned.

OpenStudy (anonymous):

3 is odd 5 is odd 8 is even

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

You are welcome